14a Category Prior Year Current Year Accounts payable 3,145.00 5,931.00 Accounts receivable 6,888.00 9,077.00 Accruals 5,620.00...

14a

Category Prior Year Current Year
Accounts payable 3,145.00 5,931.00
Accounts receivable 6,888.00 9,077.00
Accruals 5,620.00 6,076.00
Additional paid in capital 20,105.00 13,324.00
Cash ??? ???
Common Stock 2,850 2,850
COGS 22,159.00 18,327.00
Current portion long-term debt 500 500
Depreciation expense 975.00 952.00
Interest expense 1,289.00 1,136.00
Inventories 3,088.00 6,655.00
Long-term debt 16,844.00 22,324.00
Net fixed assets 75,757.00 74,071.00
Notes payable 4,001.00 6,585.00
Operating expenses (excl. depr.) 19,950 20,000
Retained earnings 35,741.00 34,774.00
Sales 46,360 45,665.00
Taxes 350 920

What is the firm's total change in cash from the prior year to the current year?

19. A stock just paid a dividend of $1.51. The dividend is expected to grow at 29.58% for three years and then grow at 3.06% thereafter. The required return on the stock is 11.74%. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

#2 A stock just paid a dividend of $1.28. The dividend is expected to grow at 26.56% for five years and then grow at 3.27% thereafter. The required return on the stock is 13.89%. What is the value of the stock?Answer format: Currency: Round to: 2 decimal places.

#3 A stock just paid a dividend of $2.78. The dividend is expected to grow at 24.80% for two years and then grow at 3.32% thereafter. The required return on the stock is 14.34%. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal places.

#4 The risk-free rate is 2.89% and the market risk premium is 4.97%. A stock with a β of 1.09 will have an expected return of ____%.

Answer format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))

#5 The risk-free rate is 1.46% and the expected return on the market 8.47%. A stock with a β of 1.29 will have an expected return of ____%.
Answer #1

19.
=1.51*(1.2958/1.1174)+1.51*(1.2958/1.1174)^2+1.51*(1.2958/1.1174)^3+1.51*(1.2958/1.1174)^3*1.0306/(11.74%-3.06%)=34.096471072187

2.
=1.28*(1.2656/1.1389)+1.28*(1.2656/1.1389)^2+1.28*(1.2656/1.1389)^3+1.28*(1.2656/1.1389)^4+1.28*(1.2656/1.1389)^5+1.28*(1.2656/1.1389)^5*1.0327/(13.89%-3.27%)=29.9722325085028

3.
=2.78*(1.248/1.1434)+2.78*(1.248/1.1434)^2+2.78*(1.248/1.1434)^2*1.0332/(14.34%-3.32%)=37.3975649406822

4.
=2.89%+1.09*4.97%=8.3073%

5.
=1.46%+1.29*(8.47%-1.46%)=10.5029%